OpenStudy (anonymous):

log4 (x2-3)+log4 10 = 1

OpenStudy (saifoo.khan):

Same base, multiply the upper terms.

OpenStudy (saifoo.khan):

The first property here, http://mathscene.co.uk/t/F/Su50k02_m04x.gif

OpenStudy (anonymous):

I know the properties but I'm not quite sure what to do once I get to log 4 10x^2 -30=1

OpenStudy (lgbasallote):

change to exponential form

OpenStudy (lgbasallote):

\[\large \log_a b = c \longrightarrow a^c = b\]

OpenStudy (saifoo.khan):

\[\Large 10x^2 -30 = 4^1\] \[\Large 10x^2 -30 = 4\]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (saifoo.khan):

Np.

OpenStudy (anonymous):

Should you solve for x?